OpenStudy (anonymous):

Find all the roots for:

OpenStudy (anonymous):

\[4x^4-32x-x^3+8 =0 \]

OpenStudy (anonymous):

You can rewrite this as:\[(4x^4 - x^2)-(32x-8)\]

OpenStudy (anonymous):

\[4x(x^3-8)-(x^3-8)=0\]

OpenStudy (anonymous):

or that too

OpenStudy (anonymous):

same result

OpenStudy (anonymous):

but, anyway, use the distributive property in reverse

OpenStudy (anonymous):

\[(4x-1)(x^3-8)=0\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[(4x-1)(x-2)(x^2+2x+4)\]

OpenStudy (anonymous):

seems about right

OpenStudy (anonymous):

okay i dont know what to do from here on, i cant seem to factor the third bracket

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well, you can't factor it into real numbers

OpenStudy (anonymous):

would i use the quadratic formula ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i get x1=\[-1+\sqrt{5}\]

OpenStudy (anonymous):

x2=

OpenStudy (anonymous):

\[-1-\sqrt{5}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

you get: \[-1 \pm i \sqrt{5}\]

OpenStudy (anonymous):

but how would i write that in brackets ?

OpenStudy (anonymous):

like if x=3 we would write x-3

OpenStudy (anonymous):

\[(x+1-i \sqrt{5})(x+1+i \sqrt{5})\]

OpenStudy (anonymous):

You essentially do the same thing, except the number you are subtracting is complex

OpenStudy (anonymous):

ohhh thanks :D

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

But i got 5's not 3 :S

OpenStudy (anonymous):

4-4(4) = -12